Given two polynomials A and B, write a program that adds the given two polynomials A and B.


Input

 5

0 -2

3 6 

4 7

1 -3

2 -1

 5

0 1

1 2

2 -4

3 3 

4 5


Your Output

12x^4 + 9x^3 - 5x^2 - 1x - 1

Expected

12x^4 + 9x^3 - 5x^2 - x - 1



Please Find Error


1
Expert's answer
2021-10-02T14:30:27-0400
import math
import numpy


def DispPoly(C,n):
    C = C[::-1]
    s = ""
    k=n-1
    for r in range(0,len(C)-2):
        if(r==0):
            if(C[r]>0): s = s + str(C[r])+"x^"+str(k)
            if(C[r]<0): s = s + str(C[r])+"x^"+str(k)
        if(r>0):
            if(C[r]>0): s = s + "+" + str(C[r])+"x^"+str(k)
            if(C[r]<0): s = s + str(C[r])+"x^"+str(k)
            
        k=k-1


    if(C[n-2]>0): s = s + "+" + str(C[n-2]) + "x"
    if(C[n-2]<0): s = s +       str(C[n-2]) + "x"


    if(C[n-1]>0): s = s + "+" + str(C[n-1])
    if(C[n-1]<0): s = s +       str(C[n-1])
    print(s)    
    
N = int(input("Enter the Order of Polynomial: "))
N = N+1
print("Enter coefficients of the two polynomiuals separated by SPACE: ")
Coeff=[]
AddP=[]
for r in range(0,N):
    s = "Enter coeff. of x^" + str(r) + str(": ")
    t = (str(input(s))).split(" ")
    temp = []
    temp.append(int(t[0]))
    temp.append(int(t[1]))
    Coeff.append(temp)
    AddP.append(sum(temp))
    
Coeff = np.array(Coeff)
Coeff=np.transpose(Coeff)
print("\nPolynomial-1:")
DispPoly(Coeff[0],N)
print("\nPolynomial-2:")
DispPoly(Coeff[1],N)
print("\nAfter Adding two polynomials: ")

DispPoly(AddP,N)